Find the slope of a line to the line that passes through the following points.
1-(4,0) and (3.8,2)
Find the slope of a line perpendicular to the line that passes through the following points.
(3,9) and (7,15)

Answer by mananth(16946)

x1 y1 x2 y2
4 0 3.8 2

slope m = (y2-y1)/(x2-x1)
( 2 - 0 )/( 3.8 - 4 )
( 2 / -0.2 )
m= -10

x1 y1 x2 y2
3 9 7 15

slope m = (y2-y1)/(x2-x1)
( 15 - 9 )/( 7 - 3 )
( 6 / 4 )
m= 3/ 2

Plug value of the slope and point( 3 , 9 ) in
Y = m x + b
9.00 = 9/2 + b
b= 9 - 9/2
b= 9/ 2
So the equation will be
Y = 3/2 x + 9/2
The slope of line perpendicular to the above line will be -2/3